Neha would retire 30 years from today and she would need ₹ 6,00,000 per year after her retirement, with the first retirement funds withdrawn one year from the day she retires. Assume a return of 7% per annum on her retirement funds and if her planning is for 25 years after retirement, Calculate:

a. How much lumpsum she should deposit in her account today so that she has enough funds for retirement?

b. How much she should deposit each year so that she has enough funds for retirement?



1
Expert's answer
2020-06-17T10:24:53-0400

a) "S=600,000*\frac{1.07^{25}-1}{0.07*1.07^{25}}=6,992,149.91"

S-sum need to start of retire

"S^*=\frac{6,992,149.91}{1.07^{30}}=918,538.58"

S*-sum which Neha need

b) "6,992,149.91=P*\frac{1.07^{30}-1}{1.07-1}"

"P=74,021.72"

P-payment every year